# Authentic Discrete Diffusion Model

### v1

@@ 3.3 Conditioning with Feature Extractors
ADD incorporates conditional signals for guidance while preserving the discrete diffusion pipeline. For image-conditioned tasks, an encoder $\mathcal{E}$ maps an input $\mathbf{x}$ to tokens:
@@
This vector $c$ is injected into the denoising network, as defined in Section 3.4. To reduce computation and prevent the learned condition from overshadowing the learning of the subsequent diffusion network, we adopt a feature reuse strategy inspired by (Li et al., 2024): (1) Expand each feature batch $K$ times, (2) Sample $K$ distinct timesteps per instance to generate noisy labels, and (3) Inject the same features into all noisy labels. This $K$-fold strategy (with $K=4$ in practice) enables multi-timestep optimization without repeated feature extraction.
@@ 3.4 Single and Multiple Token Generation
For single-label classification, ADD diffuses and denoises one-hot vectors corresponding to class labels. Conditioning (e.g., image features) guides the denoising process, while timestep-conditioned cross-entropy supervision enforces categorical fidelity. We adopt a simple Transformer to inject conditioning into the diffusion block (see Fig. 2). This setup demonstrates the stability and efficiency of ADD in the simplest categorical setting.
@@
ADD extends naturally to sequences. A sentence is represented as $\mathbf{Y}_{0}=[\mathbf{y}_{0,1},\dots,\mathbf{y}_{0,N}]$, where each $\mathbf{y}_{0,i}$ is a one-hot vector over the vocabulary. The forward process perturbs each token independently:
$q(\mathbf{y}_{t,i}\mid\mathbf{y}_{0,i})=\mathcal{N}\!\big(\mathbf{y}_{t,i};\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0,i},(1-\bar{\alpha}_{t})\mathbf{I}\big).$
@@
$p_{\theta}(\mathbf{y}_{0,i}\mid\mathbf{y}_{t,i},c)=\text{Softmax}\!\big(f_{\theta}(\mathbf{y}_{t,i},t,c)\big).$
@@
$\mathcal{L}_{\text{text}}=-\sum_{i=1}^{N}\sum_{k=1}^{K}y_{0,i}^{(k)}\log p_{\theta}(y_{0,i}^{(k)}\mid\mathbf{y}_{t,i},c).$
At inference, each token is discretized by $\arg\max$, and iterative refinement sharpens the entire sequence into a coherent sentence. Unlike autoregressive models, ADD denoises all tokens simultaneously, enabling parallel and efficient text generation. Conditioning is injected via self-attention in the diffusion block (see Fig. 2).
@@ 4.1 Experimental Setup
We evaluate ADD on two representative tasks: large-scale image classification and image-conditioned text generation. For classification, we use the ImageNet benchmark (Deng et al., 2009), following the standard data splitting protocol. Specifically, we adopt the $224\times 224$ resolution variant, which contains 1.28 million training images and 50,000 validation images across 1,000 categories. All images are normalized using channel-wise mean values of $(0.485,0.456,0.406)$ and standard deviation values of $(0.229,0.224,0.225)$. For text generation, we conduct experiments on the COCO Captions dataset (Lin et al., 2014), consisting of 82,783 training and 40,504 validation images paired with five human-annotated captions each. Images are preprocessed with standard resizing and cropping, and captions are tokenized into sequences of one-hot vectors over a vocabulary of 10,112 words.
Model Configuration.
Our full architecture contains 111 million trainable parameters. The Transformer encoder dominates the parameter budget (87M) and specializes in hierarchical feature extraction, while the diffusion module contains 24M parameters and is dedicated to discrete-space denoising. This asymmetry reflects our design principle of prioritizing feature quality while maintaining efficient generative refinement. The patchifier resolution is fixed at $224\times 224$ for image classification task and fixed at $256\times 256$ for image captioning task. Detailed layer specifications are provided in the supplementary material.
Training Protocol.
@@
ADD is trained from scratch for 500 epochs using AdamW (Loshchilov and Hutter, 2019). We adopt a base learning rate of $1\!\times\!10^{-4}$ with a weight decay of $0.3$ and effective batch size of $4096$, distributed across $8\times$ NVIDIA A100 80GB GPUs. A warmup of $20$ epochs is followed by cosine learning rate decay (Goyal et al., 2017). Gradient clipping is applied with a global norm of $3.0$, and PyTorch AMP is used for mixed precision training. For COCO captioning, we employed smaller batch sizes and weight decay values.
Infrastructure.
[…]
Evaluation Metrics.
For ImageNet classification, we report Top-1 accuracy on the validation set. For COCO captioning, our primary metric is CLIP Scores computed between image and text features; higher is better. Additionally, we also include qualitative examples to assess fluency and image–text alignment (Fig. 7).
Architecture Design.
ADD integrates a Transformer feature extractor with a diffusion-based categorical denoiser, creating a hybrid architecture optimized for representation learning and generative refinement. Crucially, the same diffusion backbone is shared across tasks: for classification, it denoises single one-hot vectors, while for captioning, it denoises token sequences in parallel.
Our multi-timestep feature reuse strategy further enhances training efficiency. For each encoder-derived embedding, we sample $K=4$ distinct diffusion timesteps per training batch, expanding supervision across multiple noise levels without recomputing features. This yields substantial savings in compute and improves convergence speed.
4.2 Single-Token Generation Task (Classification)
@@
Table 1: Comparison with state-of-the-art methods on ImageNet. All models are trained in an end-to-end manner. Backbone architectures include ViT-Base/16, ViT-Large/16, and ViT-Huge/14 (Dosovitskiy et al., 2020). Results are reported at an input resolution of $224\times 224$, with ViT-H additionally evaluated at $448\times 448$. “*” denotes results reported by (Li et al., 2024). The best result is highlighted in bold.
@@
ViT-Base (standard classification) * $\times224$ $\times16$ 87M 82.3
ViT-Large (standard classification) * $\times224$ $\times16$ 305M 82.6
ViT-Huge (standard classification) * $\times448$ $\times14$ 632M 83.1
ViT-Base (ADD with timestep-conditioned coefficients) $\times224$ $\times16$  [^1] M 82.8
ViT-Base (ADD without timestep-conditioned coefficients) $\times224$ $\times16$  [^1] M 83.0
@@ 4.2.1 Comparison with State-of-the-Art Methods
We begin by comparing our approach with state-of-the-art image classification models on ImageNet. Results are summarized in Tab. 1, from which two key observations emerge that highlight the advantages of our method.
ADD vs. Standard Classifiers.
First, we consider a direct comparison against standard classification models. To ensure fairness, we construct a counterpart model by replacing the authentic discrete diffusion module with a conventional classifier head—specifically, a linear layer that outputs classification logits—while keeping all other components identical. Concretely, this corresponds to attaching a linear classification head to the feature extractor shown in Fig. 2. The results in Table 1 are striking: ADD consistently outperforms the standard classifier baseline by a clear margin on the classification task. Our “ViT-Base (ADD)” model achieves a top-1 accuracy of 82.8%, substantially higher than its fair counterpart “ViT-Base (standard classification).” Even more remarkably, “ViT-Base (ADD)” surpasses “ViT-Large (standard classification)” variants that employ nearly three times more parameters (305M vs.  [^1] M). These comparisons strongly validate the effectiveness and efficiency of ADD.
To compare with standard classifiers methods, we implemented ViT-Base and successfully achieved an accuracy proposed in (Li et al., 2024) of 82.3%. We found that ViT-Base reached its highest accuracy around 250 epochs, after which it began to overfit and saw no further improvement. Our ADD method, however, reached the same accuracy as ViT-Base after 300 epochs and showed a trend of continued improvement. Therefore, we increased the number of ADD epochs until convergence. We finally achieved optimal performance of ADD after 500 epoch of training, 82.8%. See Fig. 3 for a detailed comparison.
Figure 3: Comparison of the accuracy of ADD and Classifier in different epochs.We trained ADD and our own state-of-the-art method on ImageNet for 300 epochs and measured the accuracy at each epoch. We found that the state-of-the-art method began to overfit after 250 epochs, while our method showed potential for further improvement even after 300 epochs. This explains why we extend the number of training epochs and ultimately achieved optimal performance (Top-1 score = 82.8) after 500 epochs.
We also tested the classification accuracy of ADD on ImageNet at different sampling steps. We found that ADD can achieve good classification results in about 10 sampling steps, and the best result is achieved after 20 iterations. Details shown as Fig. 5.
@@
Figure 5: Performance test of ADD at different sampling steps. We conducted experiments with and without timestep-conditioned coefficients respectively, and tested the classification accuracy of ADD on ImageNet at different sampling steps. We found that ADD can achieve good classification results in about 10 sampling steps, and the best result is achieved after 20 iterations. However, without the timestep-conditioned coefficients, the iterative process exhibits slightly higher accuracy
@@ 4.2.2 Ablation Analysis
We conduct rigorous ablations to examine the design choices of our framework. All variants share the same Transformer backbone (ViT-Base(Dosovitskiy et al., 2021),  [^1] M parameters) and training configuration. All controlled experiments were conducted under the condition of training for a full 500 epochs, where the model reached a converged state. The results demonstrate that by integrating several techniques, we achieve the performance presented in Table 1.
Table 2: Ablation study. We tested the accuracy of ADD using various training and generation strategies. We found that using timestep-conditioned cross entropy loss, classifier-free guidance, and “arg max + one hot” configuration achieved the best performance for ADD.
Methods Epoch Top-1 (%)
ADD (discrete diffusion with regressive loss) 400 0.13
ADD (discrete diffusion with timestep-conditioned cross entropy loss) 400 82.72
ADD (without classifier-free guidance) 500 82.36
ADD (with classifier-free guidance) 500 82.82
ADD (with timestep-conditioned coefficients) 500 82.82
ADD (without timestep-conditioned coefficients) 500 82.96
ADD (sampling with softmax) 500 82.35
ADD (sampling with $\arg\max$ + one-hot”) 500 82.82
@@
We first assess the necessity of timestep-conditioned cross-entropy alignment loss. In training diffusion models, regressive losses (e.g., MSE, $\mathcal{L}_{1}$) are typically used. In contrast, ADD employs timestep-conditioned cross-entropy to enforce prediction of $x_{0}$ (also denoted $x_{\text{start}}$). To test its importance, we replace timestep-conditioned cross-entropy with a regressive loss, training the model to predict noise as in conventional diffusion. Table 2 and Fig. 4 show that this substitution causes catastrophic degradation (82.73% $\rightarrow$ 0.13%), confirming that timestep-conditioned cross-entropy is indispensable for effective discrete-space diffusion. A comparison of the iterative processes for the two loss functions can be found in Fig. 4.
The Timestep-Conditioned Coefficient in the Cross-Entropy Loss
The effectiveness of the timestep-conditioned coefficients in enhancing the iterative process is experimentally demonstrated in Fig. 5. With the introduction of the timestep-dependent coefficient, the iterative process during generation significantly enhances the model’s ability to progressively refine its predictions, achieving remarkably high classification accuracy after sampling 10 steps. Meanwhile, it shows that removing this coefficient leads to higher accuracy but more modest improvements throughout the iterative process.
@@
Classifier-free guidance has proven crucial in continuous diffusion models for conditional generation, as it strengthens alignment between generated outputs and conditioning signals. To verify whether this property holds in our discrete setting, we remove classifier-free guidance from ADD while keeping all other components fixed. The performance drops notably (82.82% $\rightarrow$ 82.36%), as shown in Table 2. This confirms that classifier-free guidance plays an equally indispensable role in discrete diffusion, ensuring semantic consistency with conditioning inputs. Importantly, this result demonstrates that ADD retains the core mechanics of standard diffusion frameworks, thereby reinforcing our claim that ADD is an authentic discrete diffusion model.
@@
Finally, we compare two “to-one” operations used during sampling: (i) softmax-based sampling and (ii) $\arg\max$ followed by one-hot projection. As illustrated in Table 2, both strategies perform competitively, though $\arg\max$ + one-hot projection yields a slight but consistent advantage. Consequently, unless otherwise specified, all sampling results in this paper use $\arg\max$ + one-hot projection as the default strategy. For the specific accuracy changes resulting from different iteration settings during sampling, refer to Fig. 5. Futhermore, Fig. 6 demonstrates that the output distribution of our model gradually approaches a one-hot distribution as the iterations proceed.
@@
4.3 Multi-Token Generation Task (Text Generation)
Figure 7: Quality comparison of ADD and PDD. The ground-truth captions are shown alongside captions generated by ADD and PDD for representative COCO examples. ADD produces fluent and semantically accurate captions that align well with both the image content and ground-truth descriptions, whereas PDD often generates ungrammatical or semantically inconsistent text. This highlights the superiority of authentic discrete diffusion (ADD) over pseudo discrete diffusion (PDD) in multi-token text generation.

### v2

@@ 3.3 Conditioning with Feature Extractors
TDD incorporates conditional signals for guidance while preserving the discrete diffusion pipeline. For image-conditioned tasks, an encoder $\mathcal{E}$ maps an input $\mathbf{x}$ to tokens:
@@
This vector $c$ is injected into the denoising network, as defined in Section 3.4. To reduce computation and prevent the learned condition from overshadowing the learning of the subsequent diffusion network, we adopt a feature reuse strategy inspired by (li2024autoregressive): (1) Expand each feature batch $K$ times, (2) Sample $K$ distinct timesteps per instance to generate noisy labels, and (3) Inject the same features into all noisy labels. This $K$-fold strategy enables multi-timestep optimization without repeated feature extraction.
@@ 3.4 Single and Multiple Token Generation
For single-label classification, TDD diffuses and denoises one-hot vectors corresponding to class labels. Conditioning (e.g., image features) guides the denoising process, while timestep-conditioned cross-entropy supervision enforces categorical fidelity. We adopt a simple Transformer to inject conditioning into the diffusion block (see Fig. 2). This setup demonstrates the stability and efficiency of TDD in the simplest categorical setting.
@@
TDD extends naturally to sequences. A sentence is represented as $\mathbf{Y}_{0}=[\mathbf{y}_{0,1},\dots,\mathbf{y}_{0,N}]$, where each $\mathbf{y}_{0,i}$ is a one-hot vector over the vocabulary. The forward process perturbs each token independently:
$q(\mathbf{y}_{t,i}|\mathbf{y}_{0,i})=\mathcal{N}\big(\mathbf{y}_{t,i};\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0,i},(1-\bar{\alpha}_{t})\mathbf{I}\big).$
@@
$p_{\theta}(\mathbf{y}_{0,i}|\mathbf{y}_{t,i},c)=\text{Softmax}\big(f_{\theta}(\mathbf{y}_{t,i},t,c)\big).$
@@
$\mathcal{L}_{\text{text}}=-\sum_{i=1}^{N}\sum_{k=1}^{K}y_{0,i}^{(k)}\log p_{\theta}(y_{0,i}^{(k)}|\mathbf{y}_{t,i},c).$
At inference, each token is discretized by $\operatorname*{arg\,max}$, and iterative refinement sharpens the entire sequence into a coherent sentence. Unlike autoregressive models, TDD denoises all tokens simultaneously, enabling parallel and efficient text generation. Conditioning is injected via self-attention in the diffusion block (see Fig. 2).
@@ 4.1 Experimental Setup
We evaluate TDD on two representative tasks: image classification and image-conditioned text generation. The former utilized the ImageNet dataset, while the latter employed both the imagenet-1k-vl-enriched(imagenet-1k-vl-enriched) and COCO(lin2014microsoft) datasets. For classification, we use the ImageNet benchmark (5206848), following the standard data splitting protocol. Based on experimental observations on the standard ImageNet, we conducted experiments on the imagenet-1k-vl-enriched(imagenet-1k-vl-enriched) dataset to investigate the text generation performance of our method. However, given the limited prior work on this dataset, we transitioned to the COCO(lin2014microsoft) to perform more comprehensive experiments, thereby ensuring a more robust and fair evaluation of TDD’s performance.
Model and Training.
@@
TDD integrates a Transformer feature extractor with a diffusion-based categorical denoiser, creating a hybrid architecture optimized for representation learning and generative refinement. Our full architecture contains 111 million trainable parameters. The Transformer encoder dominates the parameter budget (87M) and specializes in hierarchical feature extraction, while the diffusion module contains 24M parameters and is dedicated to discrete-space denoising. TDD is trained from scratch for 500 epochs using AdamW (Loshchilov2017DecoupledWD). We adopt a base learning rate of $1\!\times\!10^{-4}$ with a weight decay of $0.3$ and effective batch size of $4096$, distributed across $8\times$ NVIDIA A100 80GB GPUs. A warmup of $20$ epochs is followed by cosine learning rate decay (Goyal2017AccurateLM). Gradient clipping is applied with a global norm of $3.0$, and PyTorch AMP is used for mixed precision training. For captioning, we employed smaller batch sizes and weight decay values.
Competitor.
We compare our method against the strongest existing discrete diffusion baseline, the mask-based discrete diffusion model (MDD). We do not include other discrete diffusion variants such as D3PM Austin2021StructuredDD, ArgMax Flows hoogeboom2021argmax, multinomial diffusion hoogeboom2021argmax, or discrete flows tran2019discrete, as prior work has consistently shown that their performance is substantially weaker than that of MDD. MDD adopts the same Transformer backbone as our model. For multi-token prediction tasks (e.g., text generation), MDD follows the standard masked-token paradigm: it randomly masks a subset of tokens and predicts the masked locations. For single-token prediction tasks, since only one token is present, MDD replaces that token with a special mask symbol and predicts the ground-truth category. This is equivalent to appending a class token to the input and using its final-layer representation to perform single-token generation (i.e., classification).
4.2 Single-Token Generation (Classification)
@@
Table 1: Comparison with state-of-the-art methods on ImageNet. All models are trained in an end-to-end manner. Backbone architectures include ViT-Base/16, ViT-Large/16, and ViT-Huge/14 (dosovitskiy2020image). Results are reported at an input resolution of $224\times 224$, with ViT-H additionally evaluated at $448\times 448$. “*” denotes results reported by (li2024autoregressive). The best result is highlighted in bold.
@@
MDD-ViT-Base (standard classification) * $224\times 224$ $16\times 16$ $87\,\mathrm{M}$ $82.3$
MDD-ViT-Large (standard classification) * $224\times 224$ $16\times 16$ $305\,\mathrm{M}$ $82.6$
MDD-ViT-Huge (standard classification) * $448\times 448$ $14\times 14$ $632\,\mathrm{M}$ $83.1$
TDD-ViT-Base (w/ timestep-conditioned coefficients) $224\times 224$ $6\times 16 11$$11\,\mathrm{M}$ $\mathbf{82.8}$
TDD-ViT-Base (w/o timestep-conditioned coefficients) $224\times 224$ $6\times 16 11$$11\,\mathrm{M}$ $\mathbf{83.0}$
@@ 4.2.1 Comparison with State-of-the-Art Methods
We begin by comparing our approach with state-of-the-art models on ImageNet. As the state-of-the-art method, MDD uses the [CLS] token as the target of diffusion, with its fully uninformative form represented like MASK. It is designed to diffuse a single label from the MASK using the MDD approach. Our results are summarized in Tab. 1, from which two key observations emerge that highlight the advantages of our method.
TDD vs. MDD.
First, we consider a direct comparison against MDD. To ensure fairness, we construct a counterpart model by replacing the TDD module with a conventional classifier head—specifically, a linear layer that outputs the token—while keeping all other components identical. Concretely, this corresponds to attaching a linear classification head to the feature extractor shown in Fig. 2. The results in Tab. 1 are striking: TDD consistently outperforms MDD by a clear margin on the classification task. Our “ViT-Base (TDD)” model achieves a top-1 accuracy of 82.8%, substantially higher than its fair counterpart “ViT-Base (standard classification).” Even more remarkably, “ViT-Base (TDD)” surpasses “ViT-Large (standard classification)” variants that employ nearly three times more parameters (305M vs.  [^1] M). These comparisons strongly validate the effectiveness and efficiency of TDD.
To compare with standard methods, we implemented ViT-Base and successfully achieved an accuracy proposed in (li2024autoregressive) of 82.3%. We found that ViT-Base reached its highest accuracy around 250 epochs, after which it began to overfit and saw no further improvement. Our TDD method, however, reached the same accuracy as ViT-Base after 300 epochs and showed a trend of continued improvement. Therefore, we increased the number of TDD epochs until convergence. We finally achieved optimal performance of TDD after 500 epoch of training, 82.8%. See Fig. 3 for a detailed comparison.
Figure 3: Comparison of TDD and MDD Accuracy Across Epochs. We trained TDD and our state-of-the-art method on ImageNet for 300 epochs, tracking accuracy at each epoch. The state-of-the-art method began to overfit after 250 epochs, whereas our method continued to improve beyond 300 epochs. This justifies extending the training to 500 epochs, ultimately achieving optimal performance (Top-1 score = 82.8).
We also tested the classification accuracy of TDD on ImageNet at different sampling steps. We found that TDD can achieve good classification results in about 10 sampling steps, and the best result is achieved after 20 iterations. Details shown as Fig. 5.
@@
Figure 5: Performance test of TDD at different sampling steps. We conducted experiments with and without timestep-conditioned coefficients respectively, and tested the classification accuracy of TDD at different sampling steps. We found that TDD can achieve strong performance in about 10 sampling steps, and the best result is achieved after 20 iterations. However, without the timestep-conditioned coefficients, the iterative process exhibits slightly higher accuracy
@@ 4.2.2 Ablation Analysis
We perform comprehensive ablations to understand the design choices behind our framework. All variants use the same Transformer backbone (ViT-Base (dosovitskiy2021image),  [^1] M parameters) and share identical training configurations. For controlled comparisons, each model is trained for a full 500 epochs to ensure convergence. These results show that incorporating several key techniques leads to the performance reported in Tab. 1.
Table 2: Ablation study. We ablate different training and generation strategies for TDD. Timestep-conditioned cross-entropy loss, classifier-free guidance, and the $\arg\max$ + one-hot” sampling scheme all lead to consistent improvements in Top-1 accuracy.
Methods Epochs Top-1 (%)
TDD (w/ regression loss) $400$ $0.13$
TDD (w/ timestep-conditioned CE loss) $400$ $\mathbf{82.72}$
TDD (w/o classifier-free guidance) $500$ $82.36$
TDD (w/ classifier-free guidance) $500$ $\mathbf{82.82}$
TDD (w/ timestep-conditioned coefficients) $500$ $82.82$
TDD (w/o timestep-conditioned coefficients) $500$ $\mathbf{82.96}$
TDD (sampling with softmax) $500$ $82.35$
TDD (sampling with “argmax + one-hot”) $500$ $\mathbf{82.82}$
@@
We first assess the necessity of timestep-conditioned cross-entropy alignment loss. In training diffusion models, regressive losses (e.g., MSE, $\mathcal{L}_{1}$) are typically used. In contrast, TDD employs timestep-conditioned cross-entropy to enforce prediction of $x_{0}$ (also denoted $x_{\text{start}}$). To test its importance, we replace timestep-conditioned cross-entropy with a regressive loss, training the model to predict noise as in conventional diffusion. Tab. 2 and Fig. 4 show that this substitution causes catastrophic degradation (82.73% $\rightarrow$ 0.13%), confirming that timestep-conditioned cross-entropy is indispensable for effective discrete-space diffusion. A comparison of the iterative processes for the two loss functions can be found in Fig. 4.
The Timestep-Conditioned Coefficient in the Cross-Entropy Loss.
The effectiveness of timestep-conditioned coefficients in improving the iterative process is shown in Fig. 5. With the timestep-dependent coefficient, the iterative process significantly enhances the model’s ability to refine predictions, achieving high classification accuracy after 10 steps. In contrast, removing the coefficient results in slightly higher accuracy but more limited improvements during iteration.
@@
Classifier-free guidance (CFG) is critical in continuous diffusion models for conditional generation. To test its relevance in our discrete setting, we removed CFG from TDD while keeping other components fixed. The performance dropped significantly (82.82% $\rightarrow$ 82.36%), as shown in Tab. 2, confirming that CFG is equally essential in discrete diffusion for ensuring semantic consistency with conditioning inputs. This result further supports our claim that TDD retains the core mechanics of standard diffusion models, affirming it as a Tweedie Discrete Diffusion model.
@@
Finally, we compare two “to-one” operations used during sampling: (i) softmax-based sampling and (ii) $\operatorname*{arg\,max}$ followed by one-hot projection. As illustrated in Tab. 2, both strategies perform competitively, though $\operatorname*{arg\,max}$ + one-hot projection yields a slight but consistent advantage. Consequently, unless otherwise specified, all sampling results in this paper use $\operatorname*{arg\,max}$ + one-hot projection as the default strategy. For the specific accuracy changes resulting from different iteration settings during sampling, refer to Fig. 5. Furthermore, Fig. 6 demonstrates that the output distribution of our model gradually approaches a one-hot distribution as the iterations proceed.
@@
4.3 Multi-Token Generation (Text Generation)
Figure 7: Quality comparison of TDD and MDD. The ground-truth captions are shown alongside captions generated by TDD and MDD for representative COCO examples. The superiority of our method over MDD is clearly demonstrated in the text comparison shown in the figure.